OpenStudy (anonymous):

what is lim x->0 (1+3x)^(1/x)

OpenStudy (anonymous):

Rather tricky, do you know about L'Hospitals Rule?

OpenStudy (anonymous):

yes but I got stuck around ln(1+3x)/ x^-1

OpenStudy (anonymous):

Should be ln(1+3x)/x

OpenStudy (anonymous):

Then when x approaches 0 you get 0/0, so you can use L'Hospitals Rule.

OpenStudy (anonymous):

because ln(1)=0

OpenStudy (anonymous):

I don"t think I did it right but is the answer 3x^2/ 1+3x

OpenStudy (anonymous):

that's not what I got, let me show you how I thought about this problem, but as I said before, it's tricky. The first thing you want to do is to get rid of that exponent, you have already done that it seems I used the following substitution \[ \large e^u = y = \ln(3x+1)^{\frac{1}{x}} \\ \therefore u = \frac{\ln(3x+1)}{x} \] now, since this substitution didn't change much we can let the same function approach zero \[ \lim_{u \to 0} = \frac{0}{0} \] next you take the derivative of the numerator and the denominator, individually, not the quotient rule, you will get \[ \large \frac{3}{3x+1}\] this limit as x approaches zero goes to 3, so u=3 and your final limit is \[ \large e^3\]

OpenStudy (anonymous):

In the first line I made a mistake, there is no natural log there yet, that's what I intended to do in the second step to get rid of the exponent.

OpenStudy (anonymous):

that makes so much more sense thanks.

OpenStudy (anonymous):

you are welcome